OpenStudy (anonymous):

3/2)x-(1/2)=y. convert to standard form. is it right to say (3/2)x-(1/2)y=c?

jimthompson5910 (jim_thompson5910):

(3/2)x-(1/2)=y 3x-1=2y ... Multiply EVERY term by the LCD 2 to clear out the fractions 3x = 2y + 1 3x -2y = 1